OpenStudy (anonymous):

An ordered pair is a solution to a system of linear equations in two variables if and only if __________. A. it satisfies the first (top) equation in the system B. it satisfies the second (bottom) equation in the system C. it satisfies at least one of the two equations in the system D. it satisfies both of the equations in the system

OpenStudy (fibonaccichick666):

What do you think?

OpenStudy (anonymous):

IDK THAT IS WHY I ASK

OpenStudy (fibonaccichick666):

Well, then let's look at an example shall we: \[x+y=2\\ 2x+3y=1\] After solving, we get x=5 and y=-3. So can you compare your answers to the answers to this system of two variables please

OpenStudy (anonymous):

I HAVE A 50% IN MATH I DON'T KNOW

OpenStudy (fibonaccichick666):

how about trying please

OpenStudy (fibonaccichick666):

I just gave you an example, where you have an ordered pair that is a solution to a system of equations of two variables. Just evaluate each answer choice and pick the best one.

OpenStudy (anonymous):

OK I WILL TRY

OpenStudy (anonymous):

But wouldn't be B?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i,ll help if x=0 in the first equation x+y=2 y would =2 right ? try that in the second equation 2x+3y=1 ( x still =0 ) 3*2=6 which is not =1 so it proves you need to satisfy both equations the last choice

OpenStudy (anonymous):

thank you @jamaica25